A 5,000 kg satellite is orbiting the earth in a ci
Answer #1

Gravitational Force = Centripetal Force
G*m*Me/r^2 = (m*v^2)/r
Substituing Values,
(6.67*10^-11 * 5.98 * 10^24)/(6.37 * 10^6 + 800 * 10^3) = v^2
v = 7460 m/s

Correct Option - (A)